# Utility-Scale Quantum State Preparation: Classical Training using Pauli Path Simulation

### v1

@@ Abstract
We use Pauli Path simulation to variationally obtain parametrized circuits for preparing ground states of various quantum many-body Hamiltonians. These include the quantum Ising model in one dimension, in two dimensions on square and heavy-hex lattices, and the Kitaev honeycomb model, all at system sizes of one hundred qubits or more, beyond the reach of exact state-vector simulation, thereby reaching utility scale. We benchmark the Pauli Path simulation results against exact ground-state energies when available, and against density-matrix renormalization group calculations otherwise, finding strong agreement. To further assess the quality of the variational states, we evaluate the magnetization in the $x$ and $z$ directions for the quantum Ising models and compute the topological entanglement entropy for the Kitaev honeycomb model. Finally, we prepare approximate ground states of the Kitaev honeycomb model with $48$ qubits, in both the gapped and gapless regimes, on Quantinuum’s System Model H2 quantum computer using parametrized circuits obtained from Pauli Path simulation. We achieve a relative energy error of approximately $5\%$ without error mitigation and demonstrate the braiding of Abelian anyons on the quantum device beyond fixed-point models.
I Introduction
@@
A natural strategy to mitigate these challenges is to offload as much computation as possible to classical machines [10, 11]. In particular, classical simulations can be used to train circuit parameters, with quantum hardware subsequently employed only for tasks such as simulating quantum dynamics from the classically trained quantum state. This hybrid approach can substantially reduce runtime and resource demands on quantum devices. However, conventional classical simulation methods face limitations: exact state-vector approaches are restricted to small system sizes, while tensor-network methods are typically effective only for shallow circuits or systems with one-dimensional geometries, with performance constrained by entanglement growth.
Complementing these methods, in this work we employ the recently developed coefficient-truncation Pauli Path simulation (PPS)—also known as sparse Pauli dynamics or Pauli propagation [12, 13, 14, 15, 16]—to simulate VQAs for ground-state preparation, targeting system sizes of around one hundred qubits, beyond the reach of exact state-vector simulation, at the so-called utility scale. Coefficient truncation is one of several truncation schemes within the PPS framework [17, 16, 18, 19, 20]. In essence, PPS estimates the expectation value of a quantum circuit by evolving observables in the Heisenberg picture, expanding them in the Pauli-string basis, and discarding terms with coefficients below a chosen threshold to reduce computational cost. This approach has already been successfully applied to quantum many-body dynamics at the utility scale [12, 13], with benchmarking against quantum hardware [2] demonstrating comparable precision in certain parameter regimes..
@@
The rest of the paper is organized as follows. In Sec. II, we provide an overview of the coefficient-truncation PPS method and its application to variational quantum algorithms. We then apply our approach to the quantum Ising model in one and two dimensions, with results presented in Sec. III and Sec. IV, respectively, using ground-state energies and magnetization as benchmarks. In Sec. V, we present results for the Kitaev honeycomb model, benchmarking both energy and topological entanglement entropy. Section VI demonstrates the implementation of parametrized circuits on quantum hardware for a 48-qubit Kitaev honeycomb system in both gapped and gapless regimes. The braiding demonstration on the quantum device highlights the topological nature of the prepared quantum states beyond fixed-point models. Finally, in Sec. VII, we summarize our findings, discuss their implications, and outline directions for future research.
II Pauli path-simulated variational quantum algorithm
@@
III Quantum Ising model in 1D
@@
In the VQA framework, the cost function and its gradient are typically evaluated on a quantum device. As described in Sec. II, we instead approximate them classically using PPS, with a default truncation threshold of $\delta_{c}=10^{-3}$. In practice, dynamically adjusting $\delta_{c}$ during optimization is advantageous. For example, in the early stages, when gradient magnitudes are relatively large, a higher threshold yields sufficiently accurate estimates while substantially reducing computational cost. Furthermore, reusing circuit parameters obtained at nearby model parameters as a warm start can significantly accelerate optimization.
@@
III.1 Transverse-field quantum Ising model
@@
III.2 Tilted-field quantum Ising model
@@
IV Quantum Ising model in 2D
@@
IV.1 Square lattice with PBC
@@
IV.2 Heavy-hex lattice with OBC
The second lattice geometry we study is the heavy-hex lattice with OBC, shown in Fig. 5(b), whose connectivity matches that of IBM quantum devices. We approximate the ground state of the quantum Ising model on this lattice using the parametrized circuit defined in Eqs. (9) and (IV), adapted to the heavy-hex connectivity, and benchmark the results against DMRG calculations with a maximum bond dimension of $\chi=500$.
@@ IV.2 Heavy-hex lattice with OBC
We emphasize that, although the ground-state properties of the quantum Ising model on both lattice geometries can be studied classically using quantum Monte Carlo [53], such methods do not produce explicit quantum circuits for approximately preparing the ground state. In contrast, our PPS parametrized circuit approach directly yields such circuits.
@@
V Kitaev honeycomb model
The last model to which we apply our method is the Kitaev honeycomb model [22], for which several works have proposed quantum circuits to prepare its ground state [54, 55, 56, 57, 58]. For convenience, we map the honeycomb lattice onto a square lattice with modified connectivity with PBC, as illustrated in Fig.9. The bonds are color coded by type: red for $x$-type bonds ($E_{X}$), green for $y$-type bonds ($E_{Y}$), and blue for $z$-type bonds ($E_{Z}$). The Hamiltonian is
@@ V Kitaev honeycomb model
VI Simulation on quantum hardware
In the previous sections, we obtained the circuit parameters using PPS. Since these parameters were obtained via a noiseless, classical approximation of the cost function, it is natural to ask how well the resulting states perform when the parameterized circuit is executed on a noisy near-term quantum device. To this end, we consider the Kitaev honeycomb model and the variational wavefunction described in Sec. V on a system of size $(N_{x},N_{y})=(8,6)$, corresponding to a total of $N=48$ qubits. We focus on two representative coupling values, $J=0.3$ and $J=0.6$, with repetition $\ell=5$. After obtaining the optimized parameters using PPS at $\delta_{c}=10^{-3}$ with respect to the proxy Hamiltonian, we prepare the PPS-trained state on the Quantinuum H2-2 quantum computer [34].
@@ VI.1 Energy
$E_{0}=-28.5876$ (20)
for $J=0.6$, yielding relative energy errors $\Delta E/|E_{0}|=4.45\%$ for $J=0.3$ and $\Delta E/|E_{0}|=6.11\%$ for $J=0.6$. These results are obtained without any error mitigation. Notably, even in the presence of hardware noise, our parameterized circuit can prepare states roughly $5\%$ in relative energy error for a system of $N=48$ qubits.
@@
$E_{\text{PPS}}=-28.3715$ (22)
for $J=0.6$. For the $J=0.3$ PPS result, we again observe the artifact that truncation can lead to a trial energy slightly below the true ground-state energy. A summary of energies obtained from the different methods is provided in Table 1.
@@ VI.2 Braiding statistics
VII Discussion
In this work, we propose using Pauli Path simulation for variational quantum algorithms in the task of ground-state preparation at the utility scale, i.e., for systems beyond the reach of exact state-vector simulation. To test this approach, we apply it to the quantum Ising model in one dimension, in two dimensions on both a square lattice with periodic boundary conditions and a heavy-hex lattice with open boundary conditions, as well as to the Kitaev honeycomb model, all at system sizes of one hundred qubits or more. Across all test cases, we benchmark Pauli Path simulation results against either exact solutions or density-matrix renormalization group results, finding remarkable agreement as quantified by the relative energy error and certain observables. For the 2D quantum Ising model on a heavy-hex lattice, the Pauli Path-simulated circuit may even outperform density-matrix renormalization group in certain parameter regimes, albeit with caveats. Finally, we validate and benchmark our approach on Quantinuum H2-2 quantum computer [34] for the Kitaev honeycomb model. For a system size of $N=48$ qubits, the prepared ground states achieve a relative energy error of roughly $5\%$ without any error mitigation. Moreover, the extraction of braiding statistics confirms that the prepared quantum states exhibit the expected topological properties.
Our Pauli Path-simulated variational quantum algorithm offers several advantages and potentials in both quantum computing application and classical algorithm. First, it provides a systematic approach for classically constructing parametrized circuits to approximately prepare the ground state of a system — a prerequisite for executing tasks such as quench dynamics, scattering simulations on a quantum machine, or probing dynamical responses. While certain models have ground states that are exactly solvable (e.g. the one-dimensional quantum Ising model and the Kitaev honeycomb model) or can be analyzed using classical numerical methods such as the density-matrix renormalization group (e.g. short-ranged interacting models in 1D) or quantum Monte Carlo methods (e.g. the two-dimensional quantum Ising model), preparing these states on a quantum machine can be challenging. The preparation procedure is not always straightforward, and even when a method exists — such as when the classical description of the wavefunction is a matrix-product state [51, 52] — state preparation can be resource-intensive. Our Pauli path-simulated variational quantum algorithm addresses these challenges by finding the quantum circuit through classical computation, which can then be executed on a quantum machine, as demonstrated in Sec. VI.
@@ VII Discussion
Improving the parameter update algorithm is another technical challenge worth to pursue. In this work, we employed SPSA combined with ADAM. While SPSA has the advantage of requiring only two evaluations of the cost function per optimization step, its stochastic nature limits the final accuracy. Exploring alternative optimization algorithms is therefore warranted. For instance, implementing automatic differentiation could reduce the number of cost function evaluations needed to compute gradients while maintaining high accuracy. Another promising option is the quantum natural gradient — a second-order derivative based optimization method that has been shown to have advantage over first-order derivative based methods such as gradient descent or ADAM [61]. Incorporating the quantum natural gradient (or an approximate version of it) into our Pauli Path–simulated variational algorithm would be an interesting direction for future work.

### v2

@@ Abstract
We use Pauli Path simulation to variationally obtain parametrized circuits for preparing ground states of various quantum many-body Hamiltonians. These include the quantum Ising model in one dimension, in two dimensions on square and heavy-hex lattices, and the Kitaev honeycomb model, all at system sizes of one hundred qubits or more — sizes at which generic quantum circuits are beyond the reach of exact state-vector simulation— thereby reaching utility scale. We benchmark the Pauli Path simulation results against exact ground-state energies when available, and against density-matrix renormalization group calculations otherwise, finding strong agreement. To further assess the quality of the variational states, we evaluate the magnetization in the $x$ and $z$ directions for the quantum Ising models and compute the topological entanglement entropy for the Kitaev honeycomb model. Finally, we prepare approximate ground states of the Kitaev honeycomb model with $48$ qubits, in both the gapped and gapless regimes, on Quantinuum’s System Model H2 quantum computer using parametrized circuits obtained from Pauli Path simulation. We achieve a relative energy error of approximately $5\%$ without error mitigation and demonstrate the braiding of Abelian anyons on the quantum device beyond fixed-point models.
1 Introduction
@@
A natural strategy to mitigate these challenges is to offload as much computation as possible to classical machines [10, 11, 12, 13]. In particular, classical simulations can be used to train circuit parameters, with quantum hardware subsequently employed only for tasks such as simulating quantum dynamics from the classically trained quantum state. This hybrid approach can substantially reduce runtime and resource demands on quantum devices. However, conventional classical simulation methods face limitations: exact state-vector approaches are restricted to small system sizes, while tensor-network methods are typically effective only for shallow circuits or systems with one-dimensional geometries, with performance constrained by entanglement growth.
Complementing these methods, in this work we employ the recently developed coefficient-truncation Pauli Path simulation (PPS)—also known as sparse Pauli dynamics or Pauli propagation [14, 15, 16, 17, 18]—to simulate VQAs for ground-state preparation, targeting system sizes of around one hundred qubits. At this scale, generic quantum circuits are beyond the reach of exact state-vector simulation, dubbed the utility scale. However, this does not preclude the possibility that other classical methods, such as tensor network methods or Monte Carlo techniques, may still be able to simulate certain problems at comparable system sizes, provided those problems exhibit favorable structure for such methods. Coefficient truncation is one of several truncation schemes within the PPS framework [19, 18, 20, 21, 22]. In essence, PPS estimates the expectation value of a quantum circuit by evolving observables in the Heisenberg picture, expanding them in the Pauli-string basis, and discarding terms with coefficients below a chosen threshold to reduce computational cost. This approach has already been successfully applied to quantum many-body dynamics at the utility scale [14, 15], with benchmarking against quantum hardware [2] demonstrating comparable precision in certain parameter regimes.
@@
The rest of the paper is organized as follows. In Sec. 2, we provide an overview of the coefficient-truncation PPS method and its application to variational quantum algorithms. We then apply our approach to the quantum Ising model in one and two dimensions, with results presented in Sec. 3 and Sec. 4, respectively, using ground-state energies and magnetization as benchmarks. In Sec. 5, we present results for the Kitaev honeycomb model, benchmarking both energy and topological entanglement entropy. Section 6 demonstrates the implementation of parametrized circuits on quantum hardware for a 48-qubit Kitaev honeycomb system in both gapped and gapless regimes. The braiding demonstration on the quantum device highlights the topological nature of the prepared quantum states beyond fixed-point models. Finally, in Sec. 7, we summarize our findings, discuss their implications, and outline directions for future research.
2 Pauli path-simulated variational quantum algorithm
@@
3 Quantum Ising model in 1D
@@
In the VQA framework, the cost function and its gradient are typically evaluated on a quantum device. As described in Sec. 2, we instead approximate them classically using PPS, with a default truncation threshold of $\delta_{c}=10^{-3}$. In practice, dynamically adjusting $\delta_{c}$ during optimization is advantageous. For example, in the early stages, when gradient magnitudes are relatively large, a higher threshold yields sufficiently accurate estimates while substantially reducing computational cost. Furthermore, reusing circuit parameters obtained at nearby model parameters as a warm start can significantly accelerate optimization.
@@
3.1 Transverse-field quantum Ising model
@@
In fact, we observe a common trend across the models examined in this work: our PPS results tend to perform less well near the critical point or in the gapless regime. We suspect that this behavior is primarily due to the design of the variational ansatz, rather than an intrinsic limitation of the PPS method itself.
In particular, we employ a class of parameterized circuits known as the Hamiltonian variational ansatz. The intuition behind this ansatz is as follows. To prepare a target state from a given initial state, where both are ground states of some Hamiltonians, one can imagine adiabatically tuning the Hamiltonian parameters, interpolating from those of the initial Hamiltonian to those of the target Hamiltonian. The Hamiltonian variational ansatz corresponds to a Trotterized version of this adiabatic evolution, with the Trotter steps promoted to variational parameters.
Heuristically, this approach performs well when the spectral gap remains open along the adiabatic path. However, its performance deteriorates, or requires a deeper circuit, when the gap shrinks or closes, as occurs near a critical point or in a gapless regime.
3.2 Tilted-field quantum Ising model
@@
4 Quantum Ising model in 2D
@@
4.1 Square lattice with PBC
@@
4.2 Heavy-hex lattice with OBC
The second lattice geometry we study is the heavy-hex lattice with OBC, shown in Fig. 6(b), whose connectivity matches that of IBM quantum devices. We approximate the ground state of the quantum Ising model on this lattice using the parametrized circuit defined in Eqs. (9) and (4), adapted to the heavy-hex connectivity, and benchmark the results against DMRG calculations with a maximum bond dimension of $\chi=500$.
@@ 4.2 Heavy-hex lattice with OBC
Third, DMRG may not be the optimal classical method for this problem. Other classical approaches, such as projected entangled pair states or quantum Monte Carlo, could potentially achieve lower trial energies. Moreover, incorporating the symmetries of the problem can further enhance the performance of these classical methods; this was not explored in the present work.
@@
We emphasize that, although the ground-state properties of the quantum Ising model on both lattice geometries can be studied classically using quantum Monte Carlo [55] or tensor-network-based methods such as projected entangled pair states, these approaches do not produce explicit quantum circuits for approximately preparing the ground state. In contrast, our PPS parametrized circuit approach directly yields such circuits.
@@
5 Kitaev honeycomb model
The last model to which we apply our method is the Kitaev honeycomb model [24], for which several works have proposed quantum circuits to prepare its ground state [56, 57, 58, 59, 60, 61]. For convenience, we map the honeycomb lattice onto a square lattice with modified connectivity with PBC, as illustrated in Fig.9. The bonds are color coded by type: red for $x$-type bonds ($E_{X}$), green for $y$-type bonds ($E_{Y}$), and blue for $z$-type bonds ($E_{Z}$). The Hamiltonian is
@@ 5 Kitaev honeycomb model
6 Simulation on quantum hardware
In the previous sections, we obtained the circuit parameters using PPS. Since these parameters were obtained via a noiseless, classical approximation of the cost function, it is natural to ask how well the resulting states perform when the parameterized circuit is executed on a noisy near-term quantum device. To this end, we consider the Kitaev honeycomb model and the variational wavefunction described in Sec. 5 on a system of size $(N_{x},N_{y})=(8,6)$, corresponding to a total of $N=48$ qubits. We focus on two representative coupling values, $J=0.3$ and $J=0.6$, with repetition $\ell=5$. After obtaining the optimized parameters using PPS at $\delta_{c}=10^{-3}$ with respect to the proxy Hamiltonian, we prepare the PPS-trained state on the Quantinuum H2-2 quantum computer [36].
@@ 6.1 Energy
$E_{0}=-28.2847$
for $J=0.6$, yielding relative energy errors $\Delta E/|E_{0}|=4.45\%$ for $J=0.3$ and $\Delta E/|E_{0}|=5.11\%$ for $J=0.6$. These results are obtained without any error mitigation. Notably, even in the presence of hardware noise, our parameterized circuit can prepare states roughly $5\%$ in relative energy error for a system of $N=48$ qubits.
@@
$E_{\text{PPS}}=-28.3715$
for $J=0.6$. We again observe the artifact that truncation can lead to a trial energy slightly below the true ground-state energy. A summary of energies obtained from the different methods is provided in Table 1.
@@ 6.2 Braiding statistics
7 Discussion
In this work, we propose using Pauli Path simulation for variational quantum algorithms in the task of ground-state preparation at the utility scale, i.e., for system sizes at which generic quantum circuits are beyond the reach of exact state-vector simulation. To test this approach, we apply it to the quantum Ising model in one dimension, in two dimensions on both a square lattice with periodic boundary conditions and a heavy-hex lattice with open boundary conditions, as well as to the Kitaev honeycomb model, all at system sizes of one hundred qubits or more. Across all test cases, we benchmark Pauli Path simulation results against either exact solutions or density-matrix renormalization group results, finding remarkable agreement as quantified by the relative energy error and certain observables. For the 2D quantum Ising model on a heavy-hex lattice, the Pauli Path-simulated circuit may even outperform density-matrix renormalization group in certain parameter regimes, albeit with caveats. Finally, we validate and benchmark our approach on Quantinuum H2-2 quantum computer [36] for the Kitaev honeycomb model. For a system size of $N=48$ qubits, the prepared ground states achieve a relative energy error of roughly $5\%$ without any error mitigation. Moreover, the extraction of braiding statistics confirms that the prepared quantum states exhibit the expected topological properties.
Our Pauli Path-simulated variational quantum algorithm offers several advantages and potentials in both quantum computing application and classical algorithm. First, it provides a systematic approach for classically constructing parametrized circuits to approximately prepare the ground state of a system — a prerequisite for executing tasks such as quench dynamics, scattering simulations on a quantum machine, or probing dynamical responses. While certain models have ground states that are exactly solvable (e.g. the one-dimensional quantum Ising model and the Kitaev honeycomb model) or can be analyzed using classical numerical methods such as the density-matrix renormalization group (e.g. short-ranged interacting models in 1D) or quantum Monte Carlo methods (e.g. the two-dimensional quantum Ising model), preparing these states on a quantum machine can be challenging. The preparation procedure is not always straightforward, and even when a method exists — such as when the classical description of the wavefunction is a matrix-product state [53, 54] — state preparation can be resource-intensive. Our Pauli path-simulated variational quantum algorithm addresses these challenges by finding the quantum circuit through classical computation, which can then be executed on a quantum machine, as demonstrated in Sec. 6.
@@ 7 Discussion
However, we caution that, due to the truncation scheme, the variational energy calculated from PPS may fall below the true ground-state energy, which could be problematic when no other reliable classical benchmark is available. In such cases, examining the scaling of the trial energy as a function of the truncation threshold can provide some confidence in the reliability of the result. Moreover, even if the PPS-computed trial energy becomes formally unphysical, the optimized parameters still define a valid trial wavefunction, which often performs well in practice. Evaluating the optimized circuit using other classical simulation methods, such as tensor-network approaches, or even executing it on a quantum device can serve as an independent validation of the variational energy.
@@
Improving the parameter update algorithm is another technical challenge worth to pursue. In this work, we employed SPSA combined with ADAM. While SPSA has the advantage of requiring only two evaluations of the cost function per optimization step, its stochastic nature limits the final accuracy. This limitation can be mitigated by lowering the learning rate and by averaging the gradient estimates over multiple evaluations. Although we did not systematically test the optimization outcomes across many different random seeds, our limited test cases suggest that different randomized initial parameters lead to consistent final results up to several significant digits.
Exploring alternative optimization algorithms is therefore warranted. For instance, implementing automatic differentiation could reduce the number of cost function evaluations needed to compute gradients while maintaining high accuracy. Another promising option is the quantum natural gradient — a second-order derivative based optimization method that has been shown to have advantage over first-order derivative based methods such as gradient descent or ADAM [64]. Incorporating the quantum natural gradient (or an approximate version of it) into our Pauli Path–simulated variational algorithm would be an interesting direction for future work.